Allen wants to buy a car that costs 17500. He already has 11,000. If he invests this at a simple interest rate of 2.3% how long will he need to invest until he has enough for the car?
Need asap!

Answer by fractalier(6550)

He will need to earn 17500 - 11000 = $6500 interest.
The formula for this is
I = Prt
so we plug in and solve for the time, t...
t = I / Pr = 6500 / (11000*.023) = 25.7 years
(I'd get a second job.)